Tohban Report 2013-05-08

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 01 May 2013
End Date: 08 May 2013
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Hazel Bain
List all reports


Solar Activity

There was a decent amount of activity this week, at least compared to recent months. There are currently 8 active regions on the disk. The most active this week was AR 11739, which spit out an M5 flare just after rotating onto the disk (on May 3rd) and many more C-flares with an occasional M in the days following. It's now in the center of the disk but has decayed significantly, so is no longer so active. I would expect that solar activity will continue at its current level (occasional C flares, nothing big).

How many GOES flares occurred?

 Flares above B, C, M, X class were      7    61     4     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      5    39     4     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      4    24     3     0

There were RHESSI flares/GOES flares 380 / 72

Memory Management

The SSR is emptying at the end of each pass set and reached a max of 27%.

Spacecraft Status

Cold plate temperatures are up slightly (due to daylight percentage), with CP2 at 111.3K. The cryocooler power is at 71W, 2.7A.

There were two commands sent to attempt to rebalance the cryocooler to a lower vibration amplitude. This commanding was done yesterday (Tuesday). Balancer amplitude and phase were changed simultaneously by 2 steps (1 step at a time). As a result, the accelerometer amplitude decreased by about 1 mg, to ~14mg. This effort will continue during the week. (See note at end.)

The spin rate is down to 14.6 rpm, so it's time for a spin up, which will probably happen today.

There was a spin rate snafu on May 3/4. Occasionally the magnetic field configuration is such that the ACS system does not sense any change in magnetic field as it spins, thinks something is wrong, and after a timeout period (5 min) goes into IDLE mode. This problem happens (usually a few occurrences close together) about once a year or so. The ACS system was commanded back to precession mode on the next available pass. No ill effects remained from this issue. Here are the sequence of events:

 2013-123-16:11 - ACS system went into IDLE mode due to spin rate miscalculation.
 2013-124-02:08 - Ground command to precession mode
 2013-124-03:42 - Autonomously transitioned back to Normal mode.

Data Gaps

There were no data gaps this week.

Detector issues

Detector 9 is once again showing occasional "dropouts" or noise, depending on how you want to think of it. These were short occurrences, not extreme, and occurred around the following times:

Since the problem is not severe, it was decided to not take any corrective action (i.e. changing the HV) yet, because it would be difficult to tell if it were helping. We'll wait for it to get worse first. In happy news, some preliminary analysis of flares during these noisy periods show *no* degradation in imaging for detector 9. One flare does show a slightly reduced flux for D9 (currently under investigation) but the other flare shows no reduced flux, meaning that the dropout-identification features in the RHESSI software are correctly identifying and correcting these dropouts. Further work and plots to follow!

Other notes

The rebalancing of the cryocooler will continue this week. The current plan is to continue changing the balancer amplitude and phase simultaneously in various steps across two axes.

Spacecraft Management

Decimation Normal/vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) Taking data at nights
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? D9 noise, slight
Personal tools